# version 0.75

### v4

@@ 10.6 Exercises
    Show that if the rank of the Kalman observability matrix is less than $d_{\mathbf{x}}$, then there are two initial states that after $d_{\mathbf{x}}$ iterations map to the same state. Hence the initial state is not uniquely determind.
@@
    Conversely, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
@@
10.4. In this exercise we conclude that the Kalman filter can be improperly learned by the class of linear predictors. Prove this fact as formally stated in Lemma 10.6. Conclud that the Kalman filter can be learned using Algorithm 6.
10.5. Prove Lemma 10.8, which is stated as follows. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
@@
    Online learning of linear predictors, which offers a relaxation of the Kalman filtering approach by directly learning predictive models without explicit system identification. Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs. While this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the system’s stability constant, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.
In this chapter, we introduce Spectral Filtering, an alternative approach that bypasses these limitations. Rather than relying on state estimation or direct system identification, spectral filtering constructs predictors in a transformed basis, which does not depend on the system’s stability properties. This method leverages spectral decompositions to efficiently extract the relevant dynamical structure, enabling more robust and scalable learning in both stable and unstable systems.
@@
While this chapter primarily focuses on symmetric system matrices for clarity, the principles of spectral filtering extend to asymmetric and more general linear dynamical system settings. These extensions require additional mathematical tools, which we reference in the bibliographic section.
@@ 11.1 Spectral filtering in One Dimension
In this section we give a simple introduction to the technique of spectral filtering, restricting ourselves to a single dimension. The restriction to a single dimension simplifies the exposition and yet is sufficient to illustrates the main ideas. In the next section we consider the high dimensional case.
The single dimensional systems which we consider next are scalar systems, with $d_{u}=d_{y}=d_{x}=1$. We assume for simplicity that the initial state is zero $\mathbf{x}_{0}=0$. Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can written and simplified to,
@@
The main idea of spectral filtering is to rewrite the vectors $\mu_{\alpha}$ in a more efficient basis, as opposed to the standard basis. The notion of efficiency stems from the definition of $\varepsilon$-approximation we have studied in previous chapters. To define the basis, we consider the following matrix:
@@ 11.1.1 The Magic of Hankel Matrices
This theorem shows that the eigenvalues of a Hankel matrix decay very rapidly. A further illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=30$ on a logarithmic scale.
@@ 11.2 Spectral Predictors
The basis of eigenvectors of a Hankel matrix gives rise to an interesting class of predictors. This class of mappings do not directly apply to the history of inputs and/or observations. Instead, we first apply a linear transformation over these vectors, and then apply a linear map. More formally,
@@
Notice that as $h\mapsto T$, this class is equivalent to the class of all linear predictors on the inputs $\prod^{LIN}$. The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what then is the benefit here?
The answer is that linear transformations are equivalent when we have full rank transforms. In our case, however, we consider low rank approximations to the transform, and the approximation properties of different subspaces can differ drastically! This is the case in our setting: the spectral basis arising from the matrix $Z_{T}$ gives excellent approximations to the class of (standard) linear predictors, as we formally prove next. The following lemma is analogous to Lemma 9.4, with the significant difference that the number of parameters $h$ does not depend on the stabilizability parameter of the system $\gamma$ at all!
@@
Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exists constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
@@
Using this property, we can prove Lemma 11.3, by simply shifting to the spectral basis representation as follows.
@@
$=\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\|$
$=\|\sum_{i=h+1}^{T}M^{\star}_{i}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}\|$
@@ 11.2.1 Generalization to High Dimension
Consider a linear dynamical system with matrices $A,B,C$ and further assume that $A$ is symmetric. This implies that $A$ is diagonalizable over the real numbers as $A=U^{\intercal}DU$. Decomposing over each dimension $j\in[d]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
@@ 11.3 Online Spectral Filtering
The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent in the online convex optimization framework, as per Algorithm 3, giving rise to the following algorithm.
@@
Figure 11.2: The filters obtained by the eigenvectors of $Z$.
The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly implies the following corollary,
@@ 11.4 Conclusion
In this chapter, we introduced Spectral Filtering as a novel approach to state estimation and prediction in linear dynamical systems, offering an alternative to classical methods like Kalman filtering and online learning of linear predictors. By leveraging spectral decompositions, we bypass the need for explicit system identification and provide a framework that is robust to system stability properties.
@@
We demonstrated the method first in the one-dimensional case, providing intuition for how spectral predictors extract relevant dynamical structure. This was then extended to higher-dimensional settings, where spectral filtering enables a structured and computationally efficient approach to forecasting. Crucially, we showed that this approach does not rely on system stability constants, making it particularly appealing for learning in both stable and unstable dynamical systems.
From a methodological standpoint, we connected Spectral Filtering to convex optimization techniques, showing that it can be implemented via online learning in a spectral basis. This formulation allows for efficient sequential learning and prediction, with provable performance guarantees.
@@ 11.5 Bibliographic Remarks
Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods.
Deviating from classical recovery and into the improper prediction world, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$. The spectral properties of Hankel matrices used in this chapter are from [BT17].
The first extension of the spectral learning technique to asymmetric matrices is due to [HLS+18]. A more recent advancement of spectral filtering to general linear dynamical systems was recently put forth in [MH25].
@@
In [AHL+18], the authors leverage a spectral filtering technique to represent linear dynamical systems concisely in a linear basis, formulating the optimal control problem as a convex program. This method circumvents the non-convex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal.
A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter in order to design neural architectures for sequence prediction. This was first suggested in the context of state space models in the spectral basis in [ASCH23], and extended to hybrid models [LND+24]. Later extensions explored efficient implementation [ACD+24] and the concept of length generalization in sequence prediction [MDA+24].
@@ 11.6 Exercises
11.3. The fact that $Z_{T}$ is positive semidefinite is important to having the eigenvalues of $Z_{T}$ decay. Give an example of a family of $n\times n$ non-positive semidefinite Hankel matrices that have eigenvalues with constant absolute value.
@@
    Find a convex set $\mathcal{M}$ of tuples $(M_{1},M_{2},\dots)$’s to optimize across such that its diameter is some $\textrm{polylog}(T,1/\varepsilon)$. The norm we measure diameter with respect to is the sum of the Frobenius norms of the $M_{i}$’s.
@@
    One can further show that the norm of the gradient is at most $\textrm{polylog}(T,1/\varepsilon)$ as well. Use these two facts to show that with the right value of $\varepsilon$, the regret against the hypothesis class $\Pi_{\star}$ is at most $\tilde{O}(\sqrt{T})$.
@@
In previous chapters, we explored control and learning techniques in fully observed linear dynamical systems, where the system state is directly accessible to the controller. However, in many real-world scenarios—such as robotics, economics, and signal processing—the state is only partially observed, requiring the controller to infer system dynamics from noisy or incomplete observations.
This chapter focuses on partially observed LDS (PO-LDS), where the state is hidden, and only a linear projection of it is observed. For the remainder of this part of the text, we restrict ourselves to time-invariant systems, and refer to the bibliographic section for extensions. The system follows the general form:
@@
Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ the control input, $\mathbf{y}_{t}$ the observed output, and $\mathbf{w}_{t}$ the disturbance. Often, partially observed LDSs are defined with an additional noise term corrupting the observations. Exercise 12.6 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
@@
In the partial observation setting some of the control policy classes from chapter 6 are not well defined. We proceed to define policy classes and describe the relationship between them.
Not all control policies designed for fully observed systems extend naturally to partially observed settings. In this chapter, we systematically explore different policy classes, analyzing their expressivity, computational tractability, and suitability for PO-LDS.
  • Linear Observation-Action Controllers (LOACs): These controllers directly map observations to actions but are highly inexpressive, as they ignore past observations and fail to capture hidden state dependencies.
  • Generalized Linear Observation-Action Controllers (GOACs): These extend LOACs by incorporating past history of observations, significantly improving expressiveness.
@@
  • Disturbance Response Controllers (DRCs): Introduced as an alternative to state-based policies, DRCs directly act on a counterfactual signal—the “Nature’s y’s”—which captures the system’s response had no control been applied. We show that DRCs can approximate LDCs, making them a powerful convex relaxation.
We conclude by analyzing the relationships between these policy classes, showing how they fit into the broader framework of nonstochastic control with partial observation. This chapter lays the groundwork for online learning and regret minimization in PO-LDS, which we explore in the next chapter.
@@ 12.1 Linear Observation Action Controllers
A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers bounded in Fronbenius norm by $\|K\|\leq\gamma$, that are also $\gamma$-stabilizing for given LDS.
While this class of policies is natural to consider, it turns out to be very inexpressive. Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 12.6). Therefore we proceed to define more expressive and useful policies. However, LOACs will be useful in simplifying proofs later on.
@@ 12.1.1 Generalized Linear Observation Action Controllers
We can generalize linear observation-action controllers to apply to a history of past observations. This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class. The formal definition is below.
@@
Denote by $\Pi^{G}_{h,\gamma}$ be that class of all such policies satisfying $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ that are also $\gamma$-stabilzing.
@@
The following lemma relates the original and lifted system (for a proof see exercise 12.6).
@@
The lemma above shows that generalized linear observation-action controllers are a special case of simple observation-action controllers such that $\mathbf{u}_{t}^{\pi}=[K_{0}\dots K_{h}]\tilde{\mathbf{y}}_{t}$ for the lifted system.
@@ 12.2 Linear Dynamic Controllers
Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers, namely GOACs. The proof is left as Exercise 12.6.

### v5

@@ 10.6 Exercises
    Show that if the rank of the Kalman observability matrix is less than $d_{\mathbf{x}}$, then there are two initial states that after $d_{\mathbf{x}}$ iterations map to the same state. Hence, the initial state is not uniquely determined.
@@
    In contrast, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
@@
10.4. In this exercise we conclude that the Kalman filter can be improperly learned by the class of linear predictors. Prove this fact as formally stated in Lemma 10.6. Conclude that the Kalman filter can be learned using Algorithm 6.
10.5. Prove Lemma 10.8, which is stated below. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
@@
    Online learning of linear predictors, which offers a relaxation of the Kalman filtering approach by directly learning predictive models without explicit system identification. Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs. Although this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the stability constant of the system, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.
In this chapter, we introduce Spectral Filtering, an alternative approach that bypasses these limitations. Rather than relying on state estimation or direct system identification, spectral filtering constructs predictors on a transformed basis, which does not depend on the stability properties of the system. This method leverages spectral decompositions to efficiently extract the relevant dynamical structure, enabling more robust and scalable learning in both stable and unstable systems.
@@
Although this chapter primarily focuses on symmetric system matrices for clarity, the principles of spectral filtering extend to asymmetric and more general linear dynamical system settings. These extensions require additional mathematical tools, which we reference in the bibliographic section.
@@ 11.1 Spectral filtering in One Dimension
In this section, we give a simple introduction to the technique of spectral filtering, restricting ourselves to a single dimension. The restriction to a single dimension simplifies the exposition and is yet sufficient to illustrate the main ideas. In the next section, we consider the high-dimensional case.
The one-dimensional systems which we consider next are scalar systems, with $d_{u}=d_{y}=d_{x}=1$. For simplicity, we assume that the initial state is zero $\mathbf{x}_{0}=0$. Recall the set of optimal predictors, the class $\Pi_{\star}$, which for scalar systems can written and simplified to,
@@
The main idea of spectral filtering is to rewrite the vectors $\mu_{\alpha}$ in a more efficient basis, as opposed to the standard basis. The notion of efficiency comes from the definition of $\varepsilon$-approximation that we have studied in previous chapters. To define the basis, we consider the following matrix:
@@ 11.1.1 The Magic of Hankel Matrices
This theorem shows that the eigenvalues of a Hankel matrix decay very rapidly. An additional illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=30$ on a logarithmic scale.
@@ 11.2 Spectral Predictors
The basis of eigenvectors of a Hankel matrix gives rise to an interesting class of predictors. This class of mappings does not directly apply to the history of inputs and/or observations. Instead, we first apply a linear transformation over these vectors, and then apply a linear map. More formally,
@@
Notice that as $h\mapsto T$, this class is equivalent to the class of all linear predictors on the inputs $\prod^{LIN}$. The reader may wonder at this point: Since a linear transformation composed with another linear transform remains linear, what is the benefit here?
The answer is that linear transformations are equivalent when we have full-rank transforms. In our case, however, we consider low-rank approximations to the transform, and the approximation properties of different subspaces can differ drastically! This is the case in our setting: the spectral basis arising from the matrix $Z_{T}$ gives excellent approximations to the class of (standard) linear predictors, as we formally prove next. The following lemma is analogous to Lemma 9.4, with the significant difference that the number of parameters $h$ does not depend on the stabilizability parameter of the system $\gamma$ at all!
@@
Let $\{\phi_{j}\}$ be the eigenvectors of Z. Then there exist constants $c_{1},c_{2}>0$ such that for all $j\in[T]$, $\alpha\in[0,1]$, we have
@@
Using this property, we can prove Lemma 11.3 by simply shifting to a representation on the spectral basis as follows.
@@
$=\left\|\sum_{i=1}^{T}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}-\sum_{i=1}^{h}M_{i}^{\star}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-i}\right\|$
$=\left\|\sum_{i=h+1}^{T}M^{\star}_{i}\phi_{i}^{\top}\tilde{\mathbf{u}}_{t-1}\right\|$
@@ 11.2.1 Generalization to High Dimension
Consider a linear dynamical system with matrices $A,B,C$ and further assume that $A$ is symmetric. This implies that $A$ can be diagonalized over the set of reals as $A=U^{\top}DU$. Decomposing over each dimension $j\in[d]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
@@ 11.3 Online Spectral Filtering
The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent (Algorithm 3) in the online convex optimization framework, which gives rise to the following algorithm.
Figure 11.2: The filters obtained by the eigenvectors of $Z$.
@@
The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly imply the following corollary.
@@ 11.4 Conclusion
In this chapter, we introduced Spectral Filtering as a novel approach to state estimation and prediction in linear dynamical systems, offering an alternative to classical methods like Kalman filtering and online learning of linear predictors. Using spectral decompositions, we bypass the need for explicit system identification and provide a framework that is robust to the stability properties of the system.
@@
We demonstrated the method first in the one-dimensional case, providing an intuition for how spectral predictors extract relevant dynamical structure. This was then extended to higher-dimensional settings, where spectral filtering enables a structured and computationally efficient approach to forecasting. Crucially, we showed that this approach does not rely on system stability constants, making it particularly appealing for learning in both stable and unstable dynamical systems.
From a methodological standpoint, we connected Spectral Filtering to convex optimization techniques, showing that it can be implemented via online learning in a spectral basis. This formulation allows for efficient sequential learning and prediction with provable performance guarantees.
@@ 11.5 Bibliographic Remarks
Deviating from classical recovery and into the world of improper prediction, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$. The methods we surveyed in previous chapters to learning linear dynamical systems had statistical and computational complexity that depended on the hidden dimension of the system. In contrast, the spectral filtering technique studied in this chapter has regret bounds and computational complexity guarantees that do not depend on the hidden dimension.
The spectral properties of the Hankel matrices used in this chapter are from [BT17].
The first extension of the spectral learning technique to asymmetric matrices is due to [HLS+18]. A more recent advancement of spectral filtering for general linear dynamical systems was recently put forth in [MH25].
@@
In [AHL+18], the authors use a spectral filtering technique to concisely represent linear dynamical systems in a linear basis, formulating the optimal control problem as a convex program. This method circumvents the nonconvex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal. Closed-loop control using spectral filtering techniques was recently advanced in [BBDH25].
A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter to design neural architectures for sequence prediction. This was first suggested in the context of state space models in the spectral basis in [ASCH23], and was extended to hybrid models in [LND+24]. Later extensions explored efficient implementation [ACD+24] and the concept of length generalization in sequence prediction [MDA+24].
@@ 11.6 Exercises
11.3. The fact that $Z_{T}$ is positive semidefinite is important to ensure that the eigenvalues of $Z_{T}$ decay. Give an example of a family of $n\times n$ non-positive semidefinite Hankel matrices that have eigenvalues with constant absolute value.
@@
    Find a convex set $\mathcal{M}$ of tuples $(M_{1},M_{2},\dots)$’s to optimize over such that its diameter is some $\textrm{polylog}(T,1/\varepsilon)$. The norm we measure diameter with respect to is the sum of the Frobenius norms of the $M_{i}$’s.
@@
    In addition, show that the norm of the gradient is at most $\textrm{polylog}(T,1/\varepsilon)$ as well. Using these two facts, show that with the right value of $\varepsilon$, the regret against the hypothesis class $\Pi_{\star}$ is at most $\tilde{O}(\sqrt{T})$.
@@
In previous chapters, we explored control and learning techniques in fully observed linear dynamical systems, where the system state is directly accessible to the controller. However, in many real-world scenarios, such as robotics, economics, and signal processing, the state is only partially observed, requiring the controller to infer system dynamics from noisy or incomplete observations.
This chapter focuses on partially observed linear dynamical systems, where the state is hidden, and only a linear projection of it is observed. For the remainder of this part of the text, we restrict ourselves to time-invariant systems and refer to the bibliographic section for extensions. The system follows the general form:
@@
Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ the control input, $\mathbf{y}_{t}$ the observed output, and $\mathbf{w}_{t}$ the disturbance. Often, partially observed LDSs are defined with an additional noise term that corrupts the observations. Exercise 12.6 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
@@
In the partial observation setting, some of the control policy classes from Chapter 6 are not well defined. We then define policy classes and describe the relationship between them.
Not all control policies designed for fully observed systems extend naturally to partially observed settings. In this chapter, we systematically explore different policy classes, analyzing their expressivity, computational tractability, and suitability for partially observed LDS.
  • Linear Observation-Action Controllers (or simply Linear Controllers): These controllers directly map observations to actions, but are highly inexpressive, as they ignore past observations and fail to capture hidden state dependencies.
  • Generalized Linear Observation-Action Controllers (or generalized linear controllers): These extend linear controllers by incorporating the history of observations, significantly improving expressiveness.
@@
  • Disturbance Response Controllers (DRCs): Introduced as an alternative to state-based policies, DRCs act directly on a counterfactual signal, the “Nature’s y’s”, which captures the response of the system that result by uniformly setting the control input to zero. We show that DRCs can approximate LDCs, indicating that they are a highly expressive and yet convex policy class.
We conclude by analyzing the relationships between these policy classes, showing how they fit into the broader framework of nonstochastic control with partial observation. This chapter lays the groundwork for online learning and regret minimization in partially observed LDS, which we will explore in the next chapter.
@@ 12.1 Linear Observation Action Controllers
A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers bounded in the Fronbenius norm by $\|K\|\leq\gamma$, which are also $\gamma$-stabilizing for given LDS.
Although this class of policies is natural to consider, it turns out to be very inexpressive. Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 12.6). Therefore, we now define more expressive and useful policies. However, linear controllers will be useful in simplifying proofs later on.
@@ 12.1.1 Generalized Linear Observation Action Controllers
We can generalize linear observation-action controllers to apply to a history of past observations. This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class. The formal definition is given below.
@@
By $\Pi^{G}_{h,\gamma}$, we denote the class of generalized linear observation-action controllers that satisfy $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ and are also $\gamma$-stabilizing.
@@
The following lemma relates the original and lifted system (for a proof, see exercise 12.6).
@@
The above lemma shows that generalized linear observation-action controllers are a special case of simple observation-action controllers such that $\mathbf{u}_{t}^{\pi}=[K_{0}\dots K_{h}]\tilde{\mathbf{y}}_{t}$ for the lifted system.
@@ 12.2 Linear Dynamic Controllers
Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers. The proof is left as Exercise 12.6.